You have 5 of your favourite books in your school bag. What is the probability that you will pull them out in alphabetical order?
1
Expert's answer
2020-04-15T10:47:56-0400

Probability without replacement


"P(alphabet)={1 \over 5}\cdot {1 \over5-1}\cdot{1 \over 5-2}\cdot{1 \over 5-3}\cdot{1 \over 5-4}={1 \over 5!}={1 \over 120}\approx0.0083"